You need $600 in 6 months for your annual liability insurance umbrella. If your investments earn 4.00% APR (compounded monthly), how much do you have to invest each month, starting next month for 4 months such that your investment will grow to just cover your insurance bill? O$154 O $151 $131 O $145 $148
Answer #1

Answer: $ 151

Rate of interest per month = 4 / 100 x 1/ 12 = 0.0033

n = 4

PVA i, n = [ { 1 - ( 1 / ( 1 + r ) n } / r ]

PVA 0.0033 , n=4 = [ { 1 - 1 / ( 1.0033 ) 4 } / 0.0033 ] = 3.96722

Monthly investment = $ 600 / 3.96722 = $ 151.24